INFRARED SPECTROSCOPY OF THE INTERACTION OF O_{2} WITH NO TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

When O_{2} trapped in solid neon at 4.3 K in sites adjacent to an NO molecule is excited by 240 to 420 nm radiation, NO_{3} is stabilized. Photoexcitation of NO_{3} to its 663-nm state, which has ^{2}E′ symmetry, leads to the re-formation of weakly interacting (O_{2})(NO) pairs trapped in solid neon. In addition to infrared absorptions near the gas-phase band centers of NO and O_{2}, the weakly bound complex shows several absorptions at lower frequencies. Experiments in which either the O_{2} or the NO moiety of the initial sample deposit is enriched in O-18 show O-isotopic randomization of the complex which results from photodestruction of the NO_{3} product.